OpenStudy (anonymous):

whats the answer to x2-9x=0 show ur work

OpenStudy (anonymous):

\[x^2 - 9x = 0\] x ( x - 9 ) = 0 x - 9 = 0 x = 9

OpenStudy (anonymous):

@destiney18

OpenStudy (anonymous):

x=0 and x=9